# Daniil Orel1, Indraneil Paul2, Iryna Gurevych1,2, and Preslav Nakov1

### v1

@@ D.1 Input Length Stress Tests
Table 10 shows that while other approaches seem to work best on short code snippets, probably because they were trained on shorter code samples (mainly functions) as shown in Table 9, our models actually get better with longer code sequences. This matters because real code is not usually just a few lines long. Another important thing is how stable our models remain across different input lengths. When we cut the input from 512 to 128 tokens, DroidDetect-Base only drops 7.28 F1-score points (from 99.18 to 91.90), and DroidDetect-Large drops just 4.34 points (from 99.25 to 94.91). This consistency suggests the generalisability of our models to various inputs.
@@ E.1 Error Analysis
As shown in Table 3, Research/DS domain consistently yields the lowest performance for both the Base and the Large DroidDetect models, likely due to the comparatively longer and more complex code typical in this domain. In Table 4, we observe that detecting machine-generated code in JavaScript is particularly challenging, which aligns with findings from Orel et al. (2025). We attribute this to the greater variability of coding practices found in JavaScript programs - modern frameworks of JavaScript mix coding paradigms (functional programming, OOP) - and it is also not typologically related to the rest of the languages considered in our work.
@@
As shown in Figure 5, for both the Base and Large DroidDetect models, most misclassifications occur between the AI-generated and Refined classes. Notably, adversarial samples are more frequently misclassified as human-written rather than as other machine-generated (or refined) classes. This suggests that our adversarial training strategy effectively makes these examples more human-like.
@@ E.2 Interpretation of Predictions
We studied the patterns behind models’ predictions, using gradient attribution Ancona et al. (2018) and the attention maps of our models. Although these results are somewhat noisy and may not always be fully faithful (Laugel et al., 2019; Lyu et al., 2024), several common patterns emerge. Notably, much of the model’s attention is directed toward punctuation symbols (e.g., colons, semicolons, arrows), consistent with the findings of Clark et al. (2019), who showed that BERT-based models frequently focus on punctuation. A more detailed gradient-based analysis reveals that stylistic choices, such as long, descriptive comments and detailed explanations, typical of LLM-generated code, significantly impact the model’s decisions. Additionally, consecutive empty lines and extra spaces influence predictions, aligning with the CatBoost feature behaviour described in Section C.2. Together, these findings suggest that the structural and stylistic conventions of LLM-generated code differ from human-written code.

### v3

@@ D.1 Input Length Stress Tests
Table˜10 shows that while other approaches seem to work best on short code snippets, probably because they were trained on shorter code samples (mainly functions) as shown in Table˜9, our models actually get better with longer code sequences. This matters because real code is not usually just a few lines long. Another important thing is how stable our models remain across different input lengths. When we cut the input from 512 to 128 tokens, DroidDetect-Base only drops 7.28 F1-score points (from 99.18 to 91.90), and DroidDetect-Large drops just 4.34 points (from 99.25 to 94.91). This consistency suggests the generalisability of our models to various inputs.
@@ E.1 Error Analysis
As shown in Table˜3, Research/DS domain consistently yields the lowest performance for both the Base and the Large DroidDetect models, likely due to the comparatively longer and more complex code typical in this domain. In Table˜4, we observe that detecting machine-generated code in JavaScript is particularly challenging, which aligns with findings from Orel et al. (2025). We attribute this to the greater variability of coding practices found in JavaScript programs - modern frameworks of JavaScript mix coding paradigms (functional programming, OOP) - and it is also not typologically related to the rest of the languages considered in our work.
@@
As shown in Figure˜5, for both the Base and Large DroidDetect models, most misclassifications occur between the AI-generated and Refined classes. Notably, adversarial samples are more frequently misclassified as human-written rather than as other machine-generated (or refined) classes. This suggests that our adversarial training strategy effectively makes these examples more human-like.
@@ E.2 Interpretation of Predictions
We studied the patterns behind models’ predictions, using gradient attribution Ancona et al. (2018) and the attention maps of our models. Although these results are somewhat noisy and may not always be fully faithful (Laugel et al., 2019; Lyu et al., 2024), several common patterns emerge. Notably, much of the model’s attention is directed toward punctuation symbols (e.g., colons, semicolons, arrows), consistent with the findings of Clark et al. (2019), who showed that BERT-based models frequently focus on punctuation. A more detailed gradient-based analysis reveals that stylistic choices, such as long, descriptive comments and detailed explanations, typical of LLM-generated code, significantly impact the model’s decisions. Additionally, consecutive empty lines and extra spaces influence predictions, aligning with the CatBoost feature behaviour described in Section˜C.2. Together, these findings suggest that the structural and stylistic conventions of LLM-generated code differ from human-written code.